# CyPortQA: Benchmarking Multimodal Large Language Models for Cyclone Preparedness in Port Operation

### v1

@@ Introduction
Existing port-preparedness tools rely on diverse meteorological data sources to estimate storm impacts on operations and determine port condition bulletins 333 The USCG issues port condition bulletins such as Whiskey, X-Ray, Yankee, and Zulu under 33 CFR § 165.781 (U.S. Government Publishing Office 2025)., which in turn regulate vessel movements and facility preparations (U.S. Coast Guard 2025). Once a tropical cyclone reaches wind-force thresholds at sea, the National Hurricane Center (NHC) begins issuing forecast products at regular intervals, including probabilistic wind-speed maps, twelve-hour wind-forecast tables, narrative advisories on expected impacts, and graphical cones of uncertainty depicting possible storm tracks (NOAA National Hurricane Center 2025a, b). Based on these products, NOAA (National Oceanic and Atmospheric Administration) and the USCG (U.S. Coast Guard) evaluate potential impacts on port operations. The USCG then applies a rule-based framework in which alerts are triggered when forecast wind speeds exceed predefined thresholds to determine preliminary operational decisions. Final port-condition bulletins are issued in consultation with port authorities. While this protocol provides a structured and actionable approach, it heavily depends on expert judgment. Additionally, other stakeholders, including vessel owners, terminal operators, and cargo interests, typically remain reactive and await official USCG notices before adjusting their operations  (Li, Zhang, and Wang 2023; Balakrishnan, Lim, and Zhang 2022).
@@
Several challenges remain in model-driven real-time decision support for port preparedness, including improving interpretation of multifaceted forecasts, enhancing estimation of potential impacts, and ensuring reliability in decision-making (U.S. Coast Guard 2025). Addressing these challenges calls for methods that can understand and reason over diverse and evloving data. Addressing these challenges requires models with robust understanding and reasoning abilities to process diverse, evolving data. Recent advances in MLLMs have demonstrated unprecedented capabilities in integrating heterogeneous inputs, such as images (Jiang et al. 2025), structured tables (Sui et al. 2024), and world knowledge (Yu et al. 2023), into coherent contextual outputs. By reasoning jointly across multiple modalities, MLLMs can capture complex relationships, adapt to incomplete or noisy information, and generate actionable insights, making them particularly promising for high-stakes applications in disaster management. To date, no study has systematically integrated dynamic tropical cyclone forecasts with ground-truth port operational impacts and official USCG port-condition bulletins. Without this integration, it is impossible to validate MLLM outputs against real-world outcomes, preventing in-depth performance assessment and slowing the development of automated decision-support tools. Hence, this study aims to address these gaps through the following contributions:
@@
Recent advances in MLLMs have demonstrated significant potential for natural hazard assessment by integrating heterogeneous data sources, such as textual reports and imagery, into a unified reasoning framework (Agarwal et al. 2020). Existing work has been increasingly applied MLLMs across three interrelated pillars in natural hazard response: situation understanding (Hughes and Clark 2025; Sun, Wang, and Peng 2023), impact assessment (Sun, Wang, and Peng 2023; Li et al. 2025), and decision making (Chen et al. 2024; Yin et al. 2024). Previous studies have focused on leveraging the multimodal reasoning and situational awareness capabilities of MLLMs to address the immediate needs of hazard detection and emergency response under the impact of earthquakes (Ma et al. 2025), hurricanes (Li et al. 2024b; Zhu et al. 2024), and wildfires (Ramesh et al. 2025; Chen et al. 2025). In addition, recent efforts have expanded the scope from a single hazard to multi-hazard environments (Zhou and Liu 2024; Wang et al. 2025a) that support a comprehensive understanding of cascading effects. Despite these capabilities, the use of MLLMs in supporting port preparedness under the threat of tropical cyclones remains largely unexplored.
@@
In this section, we describe our methodology for constructing CyPortQA. As shown in Figure 1(a), the port-preparedness workflow reacts to real-time multimodal meteorological releases to assess the evolving situation and reason about potential impacts and regulation strategies. Based on this, We sample scenarios at key decision time points, pairing each with its corresponding weather data and operational observations. All inputs are encoded into a structured JSON formate, after which QA templates automatically generate question–answer pairs. This pipeline produces a consistent, scalable benchmark for evaluating multimodal understanding and reasoning in time-critical, high-stake port preparedness contexts.
@@
NOAA tropical cyclone products provide updates every three hours as tropical cyclones move, with time-evolving cyclone information as well as meteorological conditions. As demonstrated in Figure 2, the data is organized in a time sequence following USCG evaluation time and is updated every 12 hours. At each data update interval, we extract four types of forecast products as follows:
  • 5‑day Track‑Forecast Cone with watch/warning region, a map showing the projected storm path and coastal alert zones;
  • Tropical‑Storm‑Force Wind‑Speed Probabilities Map, showing the spatial likelihood of damaging winds;
  • plain‑text Tropical Cyclone Public Advisory, summarizing position, intensity, motion, and key hazards;
  • Wind Speed Forecast Probabilities Table, a fixed‑width table listing the forecast wind‑speed probabilities at selected locations and time bins.
@@
CyPort (Kuai et al. 2025) collects 2,197 tropical‑cyclone exposures affecting U.S. principal ports from 2015 to 2023. The port operational performance is represented by the commercial‑vessel AIS (Automatic Identification System) data captured by onboard navigation‑safety devices. For each port’s exposure to a tropical cyclone, CyPort documents commercial‑vessel activity and impact metrics, including the total reduction in vessel throughput, the onset date of disruption, and the recovery date when normal operations resume. This record offers a fine-grained view of how ports react to tropical cyclones, providing ground-truth data for potential impact estimation tasks.
@@
Together, these three data sources describe real-time port operation scenarios as tropical cyclone approaching by combining real-time forecasts, observed port performance, and regulatory actions. They supply real-world test cases, complete with input meterological data and verified outputs from ports to rigorously evaluate MLLMs.
@@
Tropical cyclone information and port operational performances are sampled every twelve hours, starting 108 hours before the National Hurricane Center’s declared landfall time and ending at landfall ($T_{c}$). Each sampling step collects NOAA tropical cyclone products and port condition bulletins, as well as port operational performance (i.e., CyPort dataset). CyPort dataset collects 1,927 cyclone–port exposures from 2015 to 2023. We retain all 410 exposures in which freight through-put falls below the historical tenth percentile and select an equal number of exposures from ports within the defined impact buffer that exhibit no significant disruption to balance the classes (Kuai et al. 2025). To assess ports’ real-time preparedness responses, we expand each of the 820 exposures across every 12-hour sampling point with available multimodal weather data, yielding 2,917 scenarios. Each scenario captures a specific cyclone and port at a certain time before landfall, along with weather forecasts, port condition bulletins, and port operational impact.
We convert inputs, including meteorological data, API-crawled port bulletins, and supplementary documents, into a unified JSON format using ChatGPT o3. Domain experts then independently cross-checked a random 10% sample of scenarios and found strong alignment with the automated outputs, verifying the accuracy and reliability of the structured representations. Port performance impacts are taken directly from the structured CyPort dataset. Each JSON record specifies a cyclone identifier, a port identifier, and the number of hours before landfall, forming the metadata for our QA pipeline, Appendix 1 lists all senario encoding prompts and JOSN template.
@@
S1.1 Spatial Awareness 46507 G / T / X TF / MC / SA
S1.2 Temporal Understanding 40707 G / T / X TF / MC / SA
S1.3 Exposure Interpretation 31885 G / T / X TF / MC / SA
S1.4 Uncertainty Quantification 29066 G / T / X TF / MC / SA
S2.1 Time of Impact 5820 G + T + X SA
S2.2 Recovery Duration 2910 G + T + X SA
S2.3 Impact Severity 8730 G + T + X MC / SA
S3.1 Condition Alert 610 G + T + X MC
S3.2 Operation Planning 822 G + T + X SA
Table 1: Task breakdown in the CyPortQA benchmark. Each task targets a specific understanding or reasoning ability and is associated with a set of QA pairs grounded in multimodal inputs. Input modalities include G (Graph), T (Table), and X (Text). Question types include TF (True/False), MC (Multiple Choice), and SA (Short Answer). The symbol ‘/‘ denotes alternatives (or), while ‘+‘ indicates combination (and).
@@
The automatic QA generation process is grounded in a generalized template. For each scenario, we employ a curated set of 48 QA templates, into which scenario metadata can be systematically inserted to generate precise question–answer pairs. Each template is designed to ensure that the answer can be directly validated from the scenario encoding, thereby preserving label accuracy while evaluating the MLLM’s ability to comprehend and reason over one or more input modalities. The resulting QA pairs collectively span a wide range of understanding and reasoning tasks aligned with key stages of preparedness: situation understanding (S1), impact estimation (S2), and decision reasoning (S3).
QA generation runs only when a template’s required fields are present in a scenario’s encoded JSON. Whenever the criteria are met, the system instantiates the question–answer pair and adds it to CyPortQA. Because port-condition bulletins are not issued in most scenario, S3-type decision reasoning questions receive sparser coverage. In Table 1, CyPortQA ultimately contains 117,178 QA pairs spanning all stages of the cyclone-preparedness workflow. QA templates are shown in Appendix 2, and the design criteria for each stage are presented below.
@@
Open-ended questions require short answers, which are either numeric values (i.e., specific numbers, dates, or durations) or descriptive responses (e.g., operational suggestions). To evaluate the numeric outputs, we apply tolerance-based accuracy (Tian et al. 2025), where a prediction is considered correct if it falls within a predefined acceptable range based on domain understanding (e.g., Recovery duration is considered correct when it falls within ± 1 day of the ground truth). For descriptive answers, we adopt an LLM-as-a-judge approach, where the ChatGPT o3 is prompted with the question requirements and scoring criteria, and asked to assign a score between 0 and 1 (Chen et al. 2023; Li et al. 2024a). For S3.2, the LLM-as-a-judge is also required to classify each response as an under-reaction, normal reaction, or over-reaction. Example prompts and implementation details for the LLM-as-a-judge evaluation are provided in Appendix 3.
@@
Figure 3 shows that all MLLMs correctly identify major ports located well within the uncertainty cone, demonstrating their ability to perform coarse-grained spatial localization. However, performance degrades when ports are located on or near the edge of the uncertainty cone. For instance, only Gemma-3 correctly classifies Port of Texas City, underscoring how boundary ports confound other models. Although geographically distant from the cone, Calhoun Port are incorrectly labeled ’within’ by LLaVA-1.6, LLaMA-3.2, and Gemini-2.5. Smaller ports may have less geographical knowledge in MLLMs, and spatial-awareness questions involving them are more often misinterpreted. Taken together, these errors indicate that MLLMs still lack the fine-grained spatial understanding needed in port operations.
Temporal understanding (Figure 4) reveals another gap. Across the selected lead times (12h, 24h, 48h, and 72h to landfall), prediction accuracy tends to improve as the time approaches landfall due to the reduced forecast uncertainty in later stages. However, the smallest deviation from ground truth is reached at 24 h rather than 12h for most models (e.g., ChatGPT-4o, Gemini-2.5, and LLaVA-1.6). This indicates that even with more accurate data, the MLLM tends to predict a time that is closer to a day (i.e. 24h). This is revealed by earlier findings that MLLMs exhibit a prior distribution or bias which favors certain outputs or decisions (McCoy et al. 2024). This imprecision or “preference” can lead to inaccurate judgments in port preparedness decisions.
Figure 4: Temporal-understanding gaps. Responses to question aggregated at scenarios of 72 h, 48 h, 24 h, and 12 h before landfall. Left panel: tolerance-based accuracy; right panel: mean deviation from ground truth with 95 % confidence intervals.
@@
Figure 5 presents a scenario that reveals reasoning gaps in MLLM-generated operational responses. In this case, none of the baseline models correctly address the regulatory requirements needed for coordinated port and facility preparedness. In particular, all models fail to reference waterfront facility regulations, which are a critical part of official port readiness protocols.
These shortcomings result from two main issues: difficulty in interpreting complex or ambiguous input, and limited knowledge of domain-specific procedures. When faced with such issues, MLLMs tend to produce vague and high-level suggestions such as “review safety procedures” or “monitor conditions” without specifying actions linked to particular operational units such as general cargo vessels, ocean-going vessels, or terminal equipment. Consequently, their responses lack the precision needed to guide real-world decisions. Even the more complete responses capture only isolated elements, such as securing equipment, they still miss broader operational requirements. These results highlight the importance of grounding model outputs in accurate situational understanding and domain-specific knowledge.
@@ Conclusion
Limitations and Future Works: Although CyPortQA provides a comprehensive benchmark, it is constructed from U.S.-based tropical cyclone cases. As a result, its applicability to regions with different cyclone patterns and port operations has yet to be tested. In addition, the current LLM-as-a-judge approach relies solely on ChatGPT o3, which may introduce bias or model-specific preferences in evaluation. Future work aims to incorporate multiple LLMs as cross-judges or involve human judges to reduce evaluation bias and enhance robustness.
Data Availability
Data will be made available on request.

### v2

@@ Abstract
Code&Datasethttps://github.com/ChenchenMobility/MLLM-Bench-CyPortQA
@@ Introduction
Existing port-preparedness tools rely on diverse meteorological data sources to estimate storm impacts on operations and determine port condition bulletins  [^3] The USCG issues port condition bulletins such as Whiskey, X-Ray, Yankee, and Zulu under 33 CFR § 165.781 (U.S. Government Publishing Office 2025)., which in turn regulate vessel movements and facility preparations (USCG 2025). Once a tropical cyclone reaches wind-force thresholds at sea, the National Hurricane Center (NHC) begins issuing forecast products at regular intervals, including probabilistic wind-speed maps, twelve-hour wind-forecast tables, narrative advisories on expected impacts, and graphical cones of uncertainty depicting possible storm tracks (NOAA NHC 2025a, b). Based on these products, NOAA (National Oceanic and Atmospheric Administration) and the USCG (U.S. Coast Guard) evaluate potential impacts on port operations. The USCG then applies a rule-based framework in which alerts are triggered when forecast wind speeds exceed predefined thresholds to determine preliminary operational decisions. Final port-condition bulletins are issued in consultation with port authorities. While this protocol provides a structured and actionable approach, it heavily depends on expert judgment. Additionally, other stakeholders, including vessel owners, terminal operators, and cargo interests, typically remain reactive and await official USCG notices before adjusting their operations  (Li, Zhang, and Wang 2023; Balakrishnan, Lim, and Zhang 2022).
@@
Several challenges remain in model-driven real-time decision support for port preparedness, including improving interpretation of multifaceted forecasts, enhancing estimation of potential impacts, and ensuring reliability in decision-making (USCG 2025). Addressing these challenges calls for methods that can understand and reason over diverse and evolving data. Addressing these challenges requires models with robust understanding and reasoning abilities to process diverse, evolving data. Recent advances in MLLMs have demonstrated unprecedented capabilities in integrating heterogeneous inputs, such as images (Jiang et al. 2025), structured tables (Sui et al. 2024), and world knowledge (Yu et al. 2023), into coherent contextual outputs. By reasoning jointly across multiple modalities, MLLMs can capture complex relationships, adapt to incomplete or noisy information, and generate actionable insights, making them particularly promising for high-stakes applications in disaster management. To date, no study has systematically integrated dynamic tropical cyclone forecasts with ground-truth port operational impacts and official USCG port-condition bulletins. Without this integration, it is impossible to validate MLLM outputs against real-world outcomes, preventing in-depth performance assessment and slowing the development of automated decision-support tools. Hence, this study aims to address these gaps through the following contributions:
@@
Recent advances in MLLMs have demonstrated significant potential for natural hazard assessment by integrating heterogeneous data sources, such as textual reports and imagery, into a unified reasoning framework (Agarwal et al. 2020). Existing work has been increasingly applied MLLMs across three interrelated pillars in natural hazard response: situation understanding (Hughes and Clark 2025; Sun, Wang, and Peng 2023), impact assessment (Sun, Wang, and Peng 2023; Li et al. 2025; Zhou et al. 2021), and decision making (Chen et al. 2024; Yin et al. 2024). Previous studies have focused on leveraging the multimodal reasoning and situational awareness capabilities of MLLMs to address the immediate needs of hazard detection and emergency response under the impact of earthquakes (Ma et al. 2025), hurricanes (Li et al. 2024a; Zhu et al. 2024), and wildfires (Ramesh et al. 2025; Chen et al. 2025). In addition, recent efforts have expanded the scope from a single hazard to multi-hazard environments (Zhou and Liu 2024; Wang et al. 2025a) that support a comprehensive understanding of cascading effects. Despite these capabilities, the use of MLLMs in supporting port preparedness under the threat of tropical cyclones remains largely unexplored.
@@
This section presents the construction process of CyPortQA. As shown in Figure 1(a), the port-preparedness workflow reacts to real-time multimodal meteorological releases to assess the evolving situation and reason about potential impacts and regulation strategies. Based on this, We sample scenarios at key decision time points, pairing each with its corresponding weather data and operational observations. All inputs are encoded into a structured JSON formate, after which QA templates automatically generate question–answer pairs. This pipeline produces a consistent, scalable benchmark for evaluating multimodal understanding and reasoning in time-critical, high-stake port preparedness contexts.
@@
NOAA tropical cyclone products provide updates every three hours as tropical cyclones move, with time-evolving cyclone information as well as meteorological conditions. As demonstrated in Figure 2, the data is organized in a time sequence following USCG evaluation time and is updated every 12 hours. At each data update interval, we extract four types of forecast products: 1) 5-day Track-Forecast Cone with watch/warning region, a map showing the projected storm path and coastal alert zones; 2) Tropical-Storm-Force Wind-Speed Probabilities Map, showing the spatial likelihood of damaging winds; 3) Plain-text Tropical Cyclone Public Advisory, summarizing position, intensity, motion, cyclone watches and warnings; and 4) Wind Speed Forecast Probabilities Table, a fixed-width table listing the forecast wind-speed probabilities at selected locations and time bins.
@@
CyPort collects 2,197 tropical‑cyclone exposures affecting U.S. principal ports from 2015 to 2023 (Kuai et al. 2025). The port operational performance is represented by the commercial‑vessel AIS (Automatic Identification System) data captured by onboard navigation‑safety devices. For each port’s exposure to a tropical cyclone, CyPort documents commercial‑vessel activity and impact metrics, including the total reduction in vessel throughput, the onset date of disruption, and the recovery date when normal operations resume. This record offers a fine-grained view of how ports react to tropical cyclones, providing ground-truth data for potential impact estimation tasks.
@@
Together, these three data sources describe real-time port operation scenarios as tropical cyclone approaching by combining real-time forecasts, observed port performance, and regulatory actions. They supply real-world test cases, complete with input meteorological data and verified outputs from ports to rigorously evaluate MLLMs.
@@
Tropical cyclone information and port operational performances are sampled every twelve hours, starting 108 hours before the NHC’s declared landfall time and ending at landfall ($T_{c}$). Each sampling step collects NOAA tropical cyclone products and port condition bulletins, as well as port operational performance (i.e., CyPort dataset). CyPort dataset collects 1,927 cyclone–port exposures from 2015 to 2023. The port impact data is highly imbalanced: only about 21% of these exposures lead to port disruption where freight throughput falls below the historical tenth percentile, and 6% are accompanied by USCG Condition bulletins. To balance the classes, we retain all 410 exposures with impact on ports and select an equal number of exposures from ports within the defined impact buffer that exhibit no significant disruption to balance the classes  (Kuai et al. 2025).
To assess ports’ real-time preparedness responses, we expand each of the 820 exposures across every 12-hour sampling point with available multimodal weather data, yielding 2,917 scenarios. Each scenario captures a specific cyclone and port at a certain time before landfall, along with weather forecasts, port condition bulletins, and port operational impact.
We convert each scenario’s data, including the meteorological data, API-crawled port bulletins, and CyPort data records, into a unified JSON format using ChatGPT o3. Domain experts then independently cross-checked a random 10% sample of scenarios and found strong alignment with the automated outputs, verifying the accuracy and reliability of the structured representations. Port performance impacts are taken directly from the structured CyPort dataset. Each JSON record specifies a cyclone identifier, a port identifier, and the number of hours before landfall, forming the metadata for our QA pipeline, Appendix 1 lists all scenario encoding prompts and JOSN template.
@@
S1.1 Spatial Awareness 46507 I / T / X TF / MC / SA
S1.2 Temporal Understanding 40707 I / T / X TF / MC / SA
S1.3 Exposure Interpretation 31885 I / T / X TF / MC / SA
S1.4 Uncertainty Quantification 29066 I / T / X TF / MC / SA
S2.1 Time of Impact 5820 I + T + X SA
S2.2 Recovery Duration 2910 I + T + X SA
S2.3 Impact Severity 8730 I + T + X MC / SA
S3.1 Condition Alert 610 I + T + X MC
S3.2 Operation Planning 822 I + T + X SA
Table 1: Task breakdown in the CyPortQA benchmark. Each task targets a specific understanding or reasoning ability and is associated with a set of QA pairs grounded in multimodal inputs. Input modalities include I (Image), T (Table), and X (Text). Question types include TF (True/False), MC (Multiple Choice), and SA (Short Answer). The symbol ‘/‘ denotes alternatives (or), while ‘+‘ indicates combination (and).
@@
The automatic QA generation process is grounded in a generalized template framework. For each scenario, we employ a curated set of 48 QA templates, into which scenario metadata can be systematically inserted to generate precise QA pairs. Each template is designed to ensure that the answer can be directly validated from the scenario encoding, thereby preserving label accuracy while evaluating the MLLM’s ability to understand and reason over one or more input modalities. The resulting QA pairs collectively span a wide range of tasks aligned with key stages of preparedness: situation understanding (S1), impact estimation (S2), and decision reasoning (S3).
QA generation runs only when a template’s required fields are present in a scenario’s encoded JSON. Whenever the criteria are met, the system instantiates the QA pair and adds it to CyPortQA. Because port-condition bulletins are not issued in most scenario, S3-type decision reasoning questions receive sparser coverage. In Table 1, CyPortQA ultimately contains 117,178 QA pairs spanning all stages of the cyclone-preparedness workflow. To mitigate potential template bias, each question was paraphrased into several semantically equivalent variants for consistency testing. QA templates and paraphrasing details are provided in Appendix 2, and the design criteria for each stage are presented below.
@@
Open-ended questions require short answers, which are either numeric values (i.e., specific numbers, dates, or durations) or descriptive responses (e.g., operational suggestions). To evaluate the numeric outputs, we apply tolerance-based accuracy (Tian et al. 2025), where a prediction is considered correct if it falls within a predefined acceptable range based on domain understanding (e.g., Recovery duration is considered correct when it falls within ± 1 day of the ground truth). For descriptive answers, we adopt Multi-LLMs-as-Judge approach, where four top-ranked LLM judges (Li et al. 2024b) were used: GPT-o3, Gemini-2.5-Pro, Claude-Sonnet-4, and GPT-o4-mini. The final evaluation score of the Multi-LLM judge is computed as the average across all four models. For S3.2 task, the judges are also required to classify responses to one of the following three categories: under-reaction, proportionate reaction, or over-reaction. The final classification is determined by majority vote. Implementation details for LLM judges are provided in Appendix 3.
@@
Figure 3 shows that the MLLMs correctly identify Port Author, which is located well within the uncertainty cone, demonstrating their ability to perform coarse-grained spatial localization. However, performance degrades when ports are located on or near the edge of the uncertainty cone. For instance, only Gemma-3 correctly classifies Port of Texas City, underscoring how boundary ports confound other models. Although geographically distant from the cone, Calhoun Port are incorrectly labeled ’within’ by three models. Smaller ports may have less geographical knowledge in MLLMs, and spatial-awareness questions involving them are more often misinterpreted. Taken together, these errors indicate that MLLMs still lack the fine-grained spatial understanding in port operations.
Temporal understanding (Figure 4) reveals another gap. Across the selected lead times (12h, 24h, 48h, and 72h to landfall), prediction accuracy tends to improve as the time approaches landfall due to the reduced forecast uncertainty in later stages. However, the smallest deviation from ground truth is reached at 24 h rather than 12h for most models (e.g., ChatGPT-4o, Gemini-2.5, and LLaVA-1.6). These MLLMs tend to predict a time close to a full day (i.e. 24h). This is revealed by earlier findings that MLLMs exhibit a prior distribution or bias which favors certain outputs or decisions (McCoy et al. 2024). This imprecision or “preference” can lead to inaccurate judgments in port preparedness decisions.
Figure 4: Temporal-understanding gaps. Performance aggregated at 72, 48, 24, and 12 h before landfall. Left panel: tolerance-based accuracy; right panel: mean deviation from ground truth with 95 % confidence intervals.
@@
Figure 5 presents a scenario that reveals reasoning gaps in MLLM-generated operational responses. In this case, none of the baseline models correctly address the regulatory requirements needed for coordinated port and facility preparedness. In particular, all models fail to mention “securing waterfront facility” that form a key part of port readiness protocols.
The missing portions of the responses reveal insufficient knowledge of waterfront facility regulations, core guidance for port operations. Domain knowledge thus becomes a key bottleneck preventing MLLMs from effectively identifying the specific procedures and interdependent actions. This deficiency further results in vague, high-level responses that generalize or oversimplify operational requirements, making the operations underestimating the urgency of real-world decisions. These results highlight the importance of grounding model outputs in accurate situational understanding and domain-specific knowledge. Effective decision support requires MLLMs that integrate domain expertise to translate weather observations into precise, actionable guidance.
@@ Conclusion
Limitations and Future Works: Although CyPortQA provides a comprehensive benchmark, it is constructed from U.S.-based tropical cyclone cases. As a result, its applicability to regions with different cyclone patterns and port operations has yet to be tested. Future work should focus on expanding the dataset to include global ports and diverse hazard events, enabling cross-regional evaluation.